OpenStudy (anonymous):

2^(7x)=23 help please

OpenStudy (anonymous):

7x * ln 2 = ln 23 7x = ln 23/ ln 2 x = (ln23/ln2 ) / 7

OpenStudy (anonymous):

\[b^x=a \iff x=\frac{\ln(A)}{\ln(b)}\]